A brass plate having surface area 200 cm^2 on one side is electroplated with 0.10 mm thick silver layers on both sides using a 15 A current. Find the time taken to do the job. The specific gravity of silver is 10.5 and its atomic weight is 1.07.9 g/mol.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. I = 15 A Surface area = 200 cm^2 , Thickness = 0.1 mm Volume of Ag deposited = 200 × 0.01 = 2 cm^3 for one side For both sides, Mass of Ag = 4 × 10.5 = 42 g Z base Ag = E/F = 107.9/96500 m = ZIT ⇒ 42 = 107.9/96500 *15*T ⇒ T = 42 * 96500/107.9 * 15 = 2504.17 sec = 41.73 min = 42 min